### v1

# Sotopia-RL: Reward Design for Social Intelligence

Abstract
Social intelligence has become a critical capability for large language models (LLMs), enabling them to engage effectively in real-world social tasks such as accommodation, persuasion, collaboration, and negotiation. Reinforcement learning (RL) is a natural fit for training socially intelligent agents because it allows models to learn sophisticated strategies directly through social interactions. However, social interactions have two key characteristics that set barriers for RL training: (1) partial observability, where utterances have indirect and delayed effects that complicate credit assignment, and (2) multi-dimensionality, where behaviors such as rapport‑building or knowledge‑seeking contribute indirectly to goal achievement. These characteristics make Markov decision process (MDP)-based RL with single-dimensional episode-level rewards inefficient and unstable. To address these challenges, we propose Sotopia-RL, a novel framework that refines coarse episode-level feedback into utterance-level, multi-dimensional rewards. Utterance-level credit assignment mitigates partial observability by attributing outcomes to individual utterances, while multi-dimensional rewards capture the full richness of social interactions and reduce reward hacking. Experiments in Sotopia, an open‑ended social learning environment, demonstrate that Sotopia-RL achieves state‑of‑the‑art social goal completion scores (7.17 on Sotopia-hard and 8.31 on Sotopia-full), significantly outperforming existing approaches. Ablation studies confirm the necessity of both utterance‑level credit assignment and multi‑dimensional reward design for RL training. Our implementation is publicly available at: https://github.com/sotopia-lab/sotopia-rl.
Figure 1: Sotopia-RL enhances social intelligence across diverse tasks. Sotopia-RL outperforms Sotopia-$\pi$ in goal completion score on accommodation, persuasion, collaboration, and negotiation tasks, as evaluated by Sotopia-eval. Sotopia-$\pi$ only utilizes episode-level goal completion score as reward signals to conduct self-reinforcement, while Sotopia-RL utilizes more fine-grained reward designs for RL training.
1 Introduction
Social intelligence (Gweon et al., 2023; Mathur et al., 2024; Zhu et al., 2025) has emerged as a critical capability for large language models (LLMs), enabling them to interpret and respond effectively to human feedback in multi-turn interactions (Han et al., 2024; Yi et al., 2024). It underpins a wide range of impactful applications, including customer service (Pandya & Holia, 2023; Bamberger et al., 2023), educational tutoring (Stamper et al., 2024; Nye et al., 2023), conflict resolution (Aggrawal & Magana, 2024), and team coordination (Li et al., 2023; Guo et al., 2024). In these interactions, communication proceeds through a sequence of utterances. We define an utterance as a discrete unit of communication that the agent deliberately produces as an action to pursue its social goals. Utterances are not merely linguistic tokens; they carry intentions, emotions, and shape the trajectory of the conversation. A socially intelligent agent must therefore generate high‑quality utterances that balance two aspects of effectiveness. At the single‑step level, an utterance should be goal‑related, directly advancing the agent’s social goal. Across multiple steps, it should also promote dialogue productivity, sustaining an engaging and constructive conversation that indirectly supports long‑term goals. This raises a central question for social learning: How can we define a high‑quality utterance and enable agents to generate them effectively?
Uniqueness of social intelligence tasks. Unlike math and coding tasks commonly addressed by training LLMs with RL (Shao et al., 2024b; Wei et al., 2025; Cui et al., 2025), social tasks (like persuasion and collaboration) pose fundamentally different challenges. First, social interactions are inherently partially observable: agents can access the dialogue history but not latent factors such as intentions, emotions, or social norms that shape the outcome. As a result, social episodes are noisy, with the quality of individual utterances often only loosely correlated with the final success. In contrast, tasks like math or coding provide much clearer signals: a correct trajectory typically requires that most steps are high‑quality, with single‑step quality closely aligned to the final result. Second, social interactions are inherently multi‑dimensional: while some utterances contribute directly to goal achievement, others support it indirectly by maintaining rapport, fostering engagement, and sustaining conversational flow. Unlike math and coding tasks, where the outcome is mostly verifiable and binary (Su et al., 2025), social success must be analyzed across multiple dimensions.
Key challenges for training social agents. RL is a natural choice for social agent training because environments like Sotopia enable dynamic generation of multi‑turn social interactions, allowing agents to learn directly from feedback (Wang et al., 2024b; Ouyang et al., 2022). Unlike supervised learning, which depends on static datasets, RL captures the adaptive and evolving nature of dialogue. Effective RL training requires an online reward model (RM) to evaluate the quality of an utterance, and the unique features of social interactions make RM training difficult. Partial observability produces noisy, high‑variance episode‑level rewards, where low-quality utterances in the middle of the episode can lead to successful outcomes. We address this with utterance‑level credit assignment, which supplies fine‑grained supervision. At the same time, the multi‑dimensionality of social interactions risks RM overfitting to spurious features. To mitigate this, we incorporate additional reward dimensions—such as rapport‑building and knowledge‑seeking—that regularize learning and align the RM with broader social expectations.
Key insights. Our key insights are that, despite the complexity of social interactions, reward assignment for social interaction is both feasible and reliable when powered by LLMs: (1) Episode‑level outcomes can be decomposed into utterance‑level contributions, and LLMs can reliably infer which utterances contributed to or hindered success. Empirically, diverse LLM families (e.g., Gemini‑2.5‑flash (Google DeepMind, 2024), GPT‑4o (Hurst et al., 2024), Qwen2.5‑7b (Qwen et al., 2025)) show strong agreement in attribution scores, with Spearman correlations exceeding 0.7. (2) Evaluating utterance quality across distinct sub‑dimensions—such as rapport, engagement, and knowledge‑sharing—simplifies the credit assignment task, turning a difficult holistic judgment into a set of rubric‑based evaluations and making the reward design with less variance. Together, these insights show why utterance‑level and multi‑dimensional reward design is practical without human annotation, and why LLMs align well with human preferences.
Main discoveries. Based on these key insights, we utilize Sotopia (Zhou et al., 2023a), an open-ended social learning environment, for the RL-based training of social agents. Sotopia provides diverse social scenarios for simulation and multi-dimensional evaluation metrics named as Sotopia-eval. Experiments conducted in the Sotopia environment reveal two key findings: (1) Social agents trained with Sotopia-RL consistently outperform all baselines on social goal completion metrics evaluated by Sotopia-eval, achieving a goal completion score of 7.17 on the Sotopia-hard benchmark and 8.31 on the full Sotopia dataset. (2) Both utterance-level reward attribution and multi-dimensional reward combination for reward design are critical for stable and effective RL training in complex social scenarios. These results highlight the importance of social reward design and validate the core design principles behind Sotopia-eval —particularly the importance of evaluating social interaction quality across diverse dimensions.
2 Related Work
Social skill learning. To augment the social intelligence of agents, prior work utilizes RL to achieve this. Sotopia-$\pi$ (Wang et al., 2024b) uses self-reinforcement learning, Ndousse et al. (2021) employ conversation-level RL rewards, and Stable Alignment (Pang et al., 2024) trains via rule-based peer feedback without rewards. SDPO (Kong et al., 2025) uses preference-based tuning but overlooks utterance-level impact. Our key contribution is the design of utterance-level rewards specifically for social tasks. Instead of conducting explicit strategy injection for training (Zhang et al., 2025; Wang et al., 2025), we implicitly model social skills within the design of our rewards.
Process reward modeling. For reinforcement learning with verifiable rewards (RLVR) like math and programming, Process Reward Modelings (PRMs) have been effectively utilized. PRIME (Cui et al., 2025) assigns token-level rewards using only outcome labels, enhancing reasoning capabilities without the need for explicit process annotations. Similarly, other works (Choudhury, 2025; Wang et al., 2024a) employ Monte Carlo (MC) rollouts to compute reward targets, enabling scalable training of RL. MC methods estimate the expected value of uncertain quantities through repeated random sampling, making them particularly effective in modeling stochastic processes and complex static decision-making tasks (Barto, 2021). Designing an utterance-level reward for social tasks can also be considered as one special type of process reward. Due to the ambiguity and multi-dimensionality of social tasks, such process reward requires multi-dimensional evaluation.
Multi-objective RL. Multi-objective RL provides a foundational approach for aligning LLMs with multiple preferences by framing alignment as a multi-objective optimization problem. The most common strategy employs linear scalarization, which combines multiple reward functions into a single objective (Jang et al., 2023; Yang et al., 2024; Li et al., 2020; Zhou et al., 2023b), enabling the reuse of standard RL techniques while varying reward weights. Recent work extends this to non-linear combinations by incorporating utility functions (Cheng et al., 2025) or using LLMs to search for reward functions (Xie et al., 2024). Other studies improve reward modeling by decomposing rewards into more informative components (Mao et al., 2025; Shenfeld et al., 2025; Lee et al., 2024). Building on these directions, our work focuses on multi-dimensional reward learning for social tasks and relies on linear scalarization, where auxiliary rewards—such as relationship maintenance and knowledge seeking—are explicitly modeled to support goal completion.
3 Sotopia Environment
In this section, we formally define the Sotopia environment, covering social interactions along with their observation and action spaces, as well as social evaluations. This definition highlights the partial observability and multi‑dimensional nature of social interactions.
3.1 Social Interaction
Given the set $I=\{A,B\}$ of participants in a dialogue, social interaction can be modeled as a partially observable Markov decision process (POMDP), represented by the tuple $\left\langle\mathcal{S},\mathcal{A},\mathcal{O},T,Z,R\right\rangle$. $\mathcal{S}$ represents the set of possible social states, $\mathcal{A}$ represents the action space, $\mathcal{O}$ represents the observation space. $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function. $Z:\mathcal{S}\rightarrow\mathcal{O}$ is the observation function. $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function that reflects the overall quality of the social interaction with respect to each agent’s private goal, such as successful persuasion or mutual understanding.
Observation space. In Sotopia, the observations refer to the history of dialogue. Each agent $i\in I$ operates under partial observability, receiving at each time step a private observation $o_{t}^{i}\in\mathcal{O}_{i}$, where $\mathcal{O}_{i}\subseteq\mathcal{O}$ denotes the private observation space of agent $i$. $\mathcal{O}_{i}$ consists of all dialogue histories and contextual cues that agent $i$ can perceive, but excludes latent variables such as the partner’s private goals, beliefs, or emotions. The observation $o_{t}^{i}$ is generated by the probabilistic observation function $Z$, modeling the partial and asymmetric nature of social perception. A social episode with $T$ turns observed by agent $i$ is defined as
$\tau_{i}=\big{(}o_{0}^{i},a_{0}^{i},;o_{1}^{i},a_{1}^{i},;\ldots,;o_{T}^{i}\big{)},$ (1)
where $o_{t}^{i}\in\mathcal{O}_{i}$ is the dialogue history observed at time $t$, and $a_{t}^{i}\in\mathcal{A}_{i}$ is the utterance generated by agent $i$ at time $t$. This episode captures the full sequence of observations and actions.
Action space. An action in Sotopia is defined as an utterance, including both verbal (e.g., speaking) and non‑verbal (e.g., smiling, hugging) communication. For each agent $i$, the action space $\mathcal{A}_{i}$ contains all such communicative behaviors. At turn $t$, the agent samples $a_{t}^{i}\sim\pi_{\theta}(\cdot\mid o_{t}^{i},g_{i})$ from its policy conditioned on observation $o_{t}^{i}$ and private goal $g_{i}$. The joint actions then contribute to and become part of the next observation $o_{t+1}^{i}$.
MDP approximation. To optimize the social agent with MDP-based RL methods like GRPO (Shao et al., 2024b), we adopt an MDP approximation of the Sotopia environment. At step $t$, the state $s_{t}^{i}$ is considered as the dialogue history, together with its private social goal, the social policy $\pi_{\theta}$ outputs a distribution over utterances $a_{t}^{i}$ (i.e., $a_{t}^{i}\sim\pi_{\theta}(\cdot\mid s_{t}^{i})$), and offline rewards $r_{t}^{i}$ from Stage 1 (Figure 4) are used to train a online utterance-level reward model $R_{\psi}(s_{t}^{i},a_{t}^{i})$.
3.2 Social Evaluation
Figure 2: An example of a social task in the Sotopia environment. Tom is agent A, and Oliver is agent B. Each agent has a unique goal that is hidden from the other. “9 / 10” indicates a single-dimensional episode-level reward provided by LLMs to describe its goal completion status.
To support RL-based training, the Sotopia environment provides quantitative feedback on how well each agent achieves its goal using Sotopia-eval, a multi-dimensional evaluation suite with an LLM-based evaluator $f_{\theta}$.
Multi-dimensional evaluation.
For each agent $i$, the LLM-based evaluator produces a score conditioned on the social episode $\tau_{i}$ and the agent’s goal $g_{i}$: $G_{i}=f_{\theta}(\tau_{i},g_{i})\in\mathbb{R}$. This score represents the extent to which the agent $i$ achieves its social goal $g_{i}$ through social interactions $\tau_{i}$. In Sotopia-eval, this single score is extended to a seven-dimensional vector $\mathbf{G}_{i}\in\mathbb{R}^{7}$, expanding from goal completion (Goal) to believability (Bel), knowledge seeking (Kno), relationship maintaining (Rel), secret keeping (Sec), social rule-following (Soc), and financial/material benefits (Fin). Among these, Goal serves as the primary objective, while the others provide auxiliary signals of social interaction quality. Detailed definition for each dimension is available in Appendix §H.
4 Methodology
In this section, we introduce the details of our proposed Sotopia-RL framework. It includes two components: (1) social reward design with LLMs (2) social agent training with RL. The first part aims to provide high-quality offline reward labels for the training of utterance-level reward models, and the second part aims to use RL to optimize social agent policy.
4.1 Social Reward Design with LLMs
As discussed in Section 1, providing high‑quality utterance-level rewards for social interactions is challenging due to the features of social interactions. We simplify this reward design problem through two steps. First, rather than scoring utterances based on dialogue history, we attribute the episode‑level outcome to individual utterances based on the full episode, reducing variance in annotation. Second, we introduce a multi‑dimensional rubric that breaks down an utterance’s contribution into clearer aspects of goal achievement. Such decomposition converts a complex, noisy task into smaller, more structured judgments that LLMs can perform reliably, yielding fine‑grained and trustworthy supervision. Formally, given a social episode $\tau_{i}$, our goal is to assign each utterance $a_{t}^{i}$ an offline reward $r_{t}^{i}$ that reflects its contribution to the progression of the interaction.
Reward attribution: from episode‑level to utterance‑level. Episode‑level rewards offer only coarse supervision, since even conversations with low‑quality utterances can still lead to successful outcomes. Therefore, episode-level rewards $G_{i}$ are not an accurate estimate of utterance-level rewards for each $a_{t}^{i}$. To provide more precise feedback, we perform utterance‑level credit assignment. Advanced LLMs (e.g., GPT‑4o) evaluate each utterance within the full episode context, yielding attribution scores $\mathcal{A}_{i}(a_{t}^{i},\tau_{i})$ between 0 and 1 for each utterance $a_{t}^{i}$. Such offline attribution enables reliable credit assignment by leveraging global context. We then refine these scores using the episode‑level outcome $G_{i}$, so that strong utterances in successful episodes are emphasized, while good contributions in failed episodes remain acknowledged but proportionally down-weighted:
$r_{t}^{i}=G_{i}\cdot\mathcal{A}_{i}(a_{t}^{i},\tau_{i}).$ (2)
Figure 3: Overview of social reward design. To better describe and model the quality of an utterance in social interactions, we expand the episode-level reward (“9/10” mentioned above) from two axes: (1) expanding from episode-level into utterance-level; (2) expanding from single-dimension to multi-dimensions, expanding from goal completion (Goal) to relationship maintaining (Rel) and knowledge seeking (Kno). It allows us to have denser reward signals for RL training.
Figure 4: Overview of social agent training. Our pipeline has three stages: (1) Data preparation: Generate GPT self‑play rollouts and annotate rewards with offline inference, where an LLM evaluates and attributes rewards after the conversation ends. (2) SFT training: Train the policy and an utterance‑level reward model via fine‑tuning. (3) RL training: Continue RL self‑play with rewards from online inference, where the RM sees only the dialogue history up to the current turn for prediction.
Reward combination: from single to multi-dimension. While utterance‑level reward attribution improves granularity, attributing to a single reward $G_{i}$ alone cannot capture the quality of an utterance. High‑quality utterances not only advance the agent’s goal but also foster engagement and sustain conversational flow. To capture such broader contributions, we extend rewards beyond the primary goal completion score $G_{i}$ to additional dimensions from Sotopia-eval. Specifically, relationship maintenance (Rel) assesses whether an interaction preserves or strengthens the relationship between agents, and knowledge seeking (Kno) assesses whether two agents gain new knowledge during interactions (full definitions in Appendix §H). These two dimensions are included for attribution because they promote collaborative and information‑rich interactions and indirectly contribute to goal achievement, whereas other metrics in Sotopia-eval are often too task‑specific to generalize. Therefore, for the utterance $t$ of agent $i$, we collect attributed dimension‑wise scores $r_{t,d}^{i}$ based on Eq. 2, normalize them within each dimension, and aggregate them into a final reward:
$r_{t}^{i}=\frac{1}{N}\sum_{d=1}^{N}\gamma_{d}\cdot\frac{r_{t,d}^{i}-\min(r_{\cdot,d}^{i})}{\max(r_{\cdot,d}^{i})-\min(r_{\cdot,d}^{i})}.$ (3)
Here $N$ is the number of reward dimensions, $\gamma_{d}$ is the weight for dimension $d$, and $\min(r_{\cdot,d}),\max(r_{\cdot,d})$ denote the range of scores for that dimension across the dataset.
Overall: offline and flexible reward design. As shown in Figure 4, we begin with a single episode‑level reward and expand it along two axes to obtain denser supervision. This process yields utterance‑level multi-dimensional rewards that capture the quality of each contribution. The design is offline because it leverages the full dialogue context after the episode concludes, allowing for better evaluation without the uncertainty of real‑time inference. It is flexible because the rubric can be extended with additional dimensions and easily adapted to different social contexts—for example, prioritizing relationship building (Rel) in therapeutic dialogue or focusing on knowledge seeking (Kno) in educational tutoring.
4.2 Social Agent Training with RL
As shown in Figure 4, Stage 1 is about social reward design, and in Stages 2 and 3, our key contribution is bridging the POMDP nature of social interactions with the scalability of MDP‑based RL through RM distillation. By distilling the knowledge in offline rewards into an utterance-level RM that provides online utterance‑level feedback, we enable efficient training while teaching the policy to implicitly infer about the latent social context.
Training utterance-level reward models. Directly applying POMDPs to LLM training is impractical because belief tracking over high‑dimensional latent states in social interactions is computationally intractable, making MDP‑based RL with RM distillation the scalable alternative. As shown on the right of Figure 4, we therefore compress offline reward signals into an RM that provides utterance‑level feedback conditioned on dialogue history. While the MDP approximation alone risks overlooking latent social factors that are only partially observable in the local dialogue history $s_{t}^{i}$, our RM $R_{\psi}(s_{t}^{i},a_{t}^{i})$ is trained to recover these signals by distilling the offline rewards into a function that infers hidden context from local observations. We train $R_{\psi}$ to regress towards the offline rewards $r_{t}^{i}$ through
$\mathcal{L}_{\text{MSE}}=\mathbb{E}_{(s_{t}^{i},a_{t}^{i})}\left[\left(R_{\psi}(s_{t}^{i},a_{t}^{i})-r_{t}^{i}\right)^{2}\right].$ (4)
This allows the resulting RM $R_{\psi}$ to provide accurate, fine‑grained feedback for RL optimization within the tractable MDP framework, while still capturing the latent dimensions of social quality that a full POMDP would model.
Training policy models with single-turn online RL. As shown on the left of Figure 4, we train the social agent policy $\pi_{\theta}$ with a reward model $R_{\psi}$ that provides utterance-level feedback. Training proceeds in two stages. First, we warm up the policy with behavior cloning (BC) on GPT self-play rollouts to ensure coherent generation. We then further train the policy with GRPO (Shao et al., 2024b), adopting a single-turn formulation for efficiency. Each self-play rollout is decomposed into multiple $(s_{t}^{i},a_{t}^{i})$ pairs, and at each step $t$ the policy generates $a_{t}^{i}\sim\pi_{\theta}(\cdot\mid s_{t}^{i})$, receives a reward $R_{\psi}(s_{t}^{i},a_{t}^{i})$, and updates to maximize expected rewards. We do not add explicit reasoning traces during inference, focusing on optimizing utterances efficiently. Although simplified to single-turn updates, the distillation from offline rewards allows the agent to achieve improved performance in multi-turn interactions.
5 Experimental Settings
Model settings. We select Qwen2.5-7b-Instruct  [^1] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct as our base LLM for the training of both the policy model and reward model. We select GPT-4o for LLM-as-the-judge in Sotopia-eval.
GPT-4o as partner
Model Sotopia-all Sotopia-hard
Goal Overall Goal Overall
GPT-4o 8.19 3.76 6.97 3.46
Claude-Sonnet-3.5 8.42 3.77 6.64 3.30
Deepseek-v3 8.14 3.72 6.69 3.31
Qwen2.5-7B +PPDPP 8.07 3.71 6.76 3.35
+EPO 8.41 3.86 6.81 3.51
+DAT 8.11 3.70 6.78 3.36
+DSI 8.15 3.70 6.87 3.42
+Sotopia-RL 8.31 3.90 7.17 3.61
Table 1: Our method outperforms state-of-the-art models when choosing GPT-4o as partner ($p<0.05$, paired t-test on the Goal dimension). Qwen2.5-7B refers to Qwen-2.5-7B-Instruct. Training method baselines from PPDPP to DSI have details available in Section §5. Full experimental results are available in Appendix §E.1.
Behavior Cloning (BC) model as partner
Attribution Dimension Sotopia-hard
Rel Kno Goal Overall
Qwen2.5-7B Uniform Goal 1.84 4.14 5.61 2.95
Singular Goal 2.74 4.93 6.64 3.41
Scaled Goal 1.82 3.83 6.74 3.15
Direct Rel 3.61 4.92 7.24 3.60
Direct Kno 2.56 6.06 6.93 3.61
Direct Goal 2.49 4.94 7.21 3.49
+Behavior Cloning (BC) 2.49 3.37 6.76 3.16
+Sotopia-$\pi$ 2.41 3.66 6.84 3.20
+Sotopia-RL 3.41 5.53 7.81 3.80
Table 2: Our social reward designs outperform reward design baselines with the BC model as partner ($p<0.05$, paired t-test on the Goal dimension). All results use Qwen2.5-7B-Instruct as the base model. Goal-RL refers to direct+Goal; Sotopia-RL combines Rel, Kno, and Goal via direct attribution by averaging. BC and Sotopia-$\pi$ are baselines without reward design. Full experimental results are provided in Appendix §E.1.
Evaluation settings. We evaluate our method on two configurations of the Sotopia benchmark: (1) Sotopia-hard, and (2) Sotopia-all. Sotopia-hard is a subset of Sotopia-all, consisting of 14 challenging social scenarios identified as difficult among all scenarios, and we use 10 distinct agent pairings per scenario. For Sotopia-all, we evaluate on the full coverage of 90 social scenarios, using 2 agent combos per scenario to ensure diversity while maintaining scalability. More statistical information is in Appendix §A.
Training method baselines. To compare the effectiveness of our training methods, we include (1) behavior cloning (BC) that utilizes social interaction trajectories between GPT-4o, which is the same as Sotopia-$\pi$; (2) Sotopia-$\pi$ (Wang et al., 2024b) that utilizes behavior cloning and self-reinforcement; (3) other most recent baselines: PPDPP (Deng et al., 2024), EPO (Liu et al., 2025), DAT (Li et al., 2024), and DSI (Zhang et al., 2025). Sotopia-RL denotes our proposed approach, which combines direct utterance-level attribution with a multi-dimensional reward combination design (Rel +Kno +Goal), trained using single-turn online RL (GRPO) without explicit reasoning but just utterance generation. Training details are available in Appendix §B.
Reward attribution baselines. To assess the effectiveness of our reward attribution strategy, we compare it against four baselines, each defining how utterance-level rewards $r_{t}^{i}$ are derived from the episode-level score $G_{i}$: (1) Uniform — every utterance receives the same reward, $r_{t}^{i}=G_{i}$ for all $t$; (2) Singular — only one selected utterance $a_{k}^{i}$ is assigned the full reward, $r_{t}^{i}=G_{i}$ if $t=k$ and $r_{t}^{i}=0$ otherwise; (3) Scaled — the episode-level reward is distributed proportionally, $r_{t}^{i}=\alpha_{t}G_{i}$ with $\sum_{t}\alpha_{t}=1$ and $\alpha_{t}\geq 0$; and (4) Direct — each utterance is independently attributed with a normalized weight, $r_{t}^{i}=\alpha_{t}G_{i}$, where each $\alpha_{t}\in[0,1]$ reflects its contextual attribution score, ensuring no utterance exceeds the episode-level score and same with Eq. 2. Direct attribution is the method used in Sotopia-RL. Details for each attribution method are in Appendix §G.
Reward combination baselines. We train our RM with labels combining three Sotopia-eval dimensions: relationship maintenance (Rel), knowledge seeking (Kno), and goal completion (Goal). Formally, the single-dimension baselines use $r_{t}^{i}=r_{t,\textsc{Rel}}^{i}$, $r_{t}^{i}=r_{t,\textsc{Kno}}^{i}$, or $r_{t}^{i}=r_{t,\textsc{Goal}}^{i}$, corresponding to Rel-only, Kno-only, and Goal-only for Eq. 3, respectively. Sotopia-RL averages the three signals, $r_{t}^{i}=\tfrac{1}{3}\big{(}r_{t,\textsc{Rel}}^{i}+r_{t,\textsc{Kno}}^{i}+r_{t,\textsc{Goal}}^{i}\big{)}$ for Eq. 3, allowing us to isolate the contribution of each component while validating the benefit of multi-dimensional reward modeling. We use a simple average to give equal importance to each reward dimension. Conducting a more fine‑grained weighting would require hyper‑parameter search, which is computationally expensive in the Sotopia environment. See Appendix §H for details.
6 Experimental Results
Sotopia-RL helps build the state-of-the-art social agents on the Sotopia benchmark. In Table 1, Qwen-2.5-7B-Instruct trained with Sotopia-RL reaches the highest goal completion score, achieving the 7.17 score in the Sotopia-hard. It indicates that our utterance-level RM provides better guidance during the training of RL. It also indicates that for multi-turn social interactions, improving the quality of single-turn interactions with suitable single-turn rewards can effectively optimize multi-turn performance. Notably, AMPO (Wang et al., 2025) reaches 7.50 on Sotopia-hard. But it includes an explicit reasoning process and requires more than 640 inference tokens per utterance on average. Therefore, it is unfair to compare AMPO with ours since we only utilize GRPO to generate utterances without extra tokens for reasoning. Full results are available in Appendix §E.
Sotopia-RL goes beyond distillation from GPT. Our training pipeline begins with GPT-based self-play episodes and GPT-based offline reward annotations. Importantly, GPT annotations are applied offline, conditioning on the entire episode, whereas during RL training, rewards are computed online, conditioned only on the preceding dialogue history. As shown in Table 1, Sotopia-RL not only matches but surpasses GPT-4o when used directly as a policy model (7.17 vs. 6.97). If Sotopia-RL were merely a stronger form of distillation, as in behavior cloning, it could at best equal GPT-4o’s performance, not exceed it.
Figure 5: Evaluation results with different LLM-based evaluators. The consistent improvement on evaluators indicates no reward hacking. Full results in Appendix §E.2.
Figure 6: Evaluation results with different partner models. The consistent improvement with multiple partners indicates no reward hacking. Full results in Appendix §E.2.
Figure 7: Goal score curve during the training process. Incorporating additional rewards into training delays convergence compared to using the Goal reward alone.
Reward attribution contributes to the performance improvement. Based on Table 2, we find that compared with different baseline methods for reward attribution, our proposed reward attribution methods (direct) bring the most significant improvement in goal completion dimensions, increasing goal completion score from 6.74 to 7.21. Our attribution methods have a performance that is much higher than uniform baselines, indicating that the attributed fine-grained dense rewards play an important role during RL training. Moreover, compared with baselines such as scaled and singular attribution, we find that relaxing attribution constraints allows LLMs greater freedom to assign scores within minimal range limits, better leveraging their social reasoning abilities and leading to superior performance.
Reward combination contributes to the performance improvement. Based on Table 2, training with multiple reward dimensions—goal completion (Goal), relationship maintenance (Rel), and knowledge seeking (Kno)—significantly improves performance compared to using a single reward dimension. The best result comes from combining all three dimensions, yielding a 7.9This improvement arises because multi‑objective RL encourages the policy to balance different aspects of interaction quality when generating utterances. Notably, as shown in Table 2, optimizing each dimension independently also improves performance on the others, suggesting that the objectives are correlated. Thus, combining them makes reward model training more robust.
7 Discussion
To assess the effectiveness of Sotopia-RL, we first ensure that its performance gains are genuine and not the result of reward hacking (RQ1). We then analyze how our improvements come from the design of the reward attribution (RQ2) and the reward combination (RQ3).
RQ1: Does our improvement come from reward hacking or shortcut learning? No, Sotopia-RL learns high-quality social skills instead of overfitting on partner models or evaluator models.
Reward hacking occurs when performance improvements are confined to a specific partner model, a particular evaluator, or fail to generalize to human interactions. In Figure 7 and Figure 7, we conduct a thorough analysis and show that the performance gains of Sotopia-RL are consistent across settings. Specifically, the improvements hold when switching between five different partner models and five different evaluator models, demonstrating strong robustness. Moreover, Table 3 confirms that these gains extend to human evaluation, further validating that the improvements are not evaluator-specific artifacts. Finally, Appendix §E.4 and §E.5 provide additional evidence from our safety and diversity evaluations. These results show that our trained policy model does not exhibit shortcut degeneration and remains safe and diverse.
Table 3: Human evaluation results for Sotopia-RL. Sotopia-RL has higher goal completion scores than other baselines for human evaluations. GPT-4o and human beings are separately used as evaluators for human evaluation. More details about human evaluation are available in Appendix §C.
Model GPT-4o Human Correlation
Sotopia-$\pi$ 6.84 5.41 0.674
Goal-RL 7.21 5.80 0.754
Sotopia-RL 7.81 5.89 0.866
Table 4: Ablation study on the method of reward attribution. We compare online reward labels, assigned by LLMs during the course of a conversation, with offline reward labels, assigned by LLMs after the conversation concludes. In both settings, the RMs and policies are trained under the same settings.
Training Method Goal Overall
Behavior Cloning (BC) 6.76 3.16
RL w/ online reward labels 6.69 3.15
RL w/ offline reward labels 7.81 3.80
RQ2: Why does utterance-level reward attribution bring improvement? The key to effective reward design lies in offline attribution, rather than in using a strong LLM.
Unlike standard MDP tasks, social interactions cannot be accurately evaluated based only on the preceding dialogue context—the quality of an utterance often depends on how the entire conversation unfolds. To address this, we attribute episode-level rewards to each utterance using information from the whole dialogue, making the reward attribution inherently offline. We point out that offline attribution is the key to our improvement. Table 4 compares two settings for training utterance-level reward models: (1) online reward labels attributed using only the preceding dialogue history, and (2) offline reward labels attributed offline using the full episode. The offline approach achieves a substantially higher goal score (7.81) than the online approach, clearly demonstrating its effectiveness. Importantly, such an improvement does not rely on GPT-4o itself. As shown in Figure 8, replacing GPT-4o with weaker models for utterance-level reward labeling produces highly correlated reward signals (¿0.7). This indicates that with well-designed prompts, precise offline credit assignment for utterance-level rewards can be reliably achieved even without state-of-the-art LLMs. More in-depth analysis and human evaluation results on utterance-level rewards are available in Appendix §F.
Figure 8: Pairwise reward label correlation. Reward labels with various LLMs are highly correlated.
RQ3: Why does the reward combination bring improvement? Using rewards with multiple dimensions makes RM training more robust, and a better RM helps prevent RL from overfitting.
To discuss why the reward combination brings improvement, we first rule out the possibility that the observed gains are merely due to reward label smoothing. To test this, we increased the attribution granularity from a 3‑point to a 10‑point scale and reran the pipeline. The 10‑point scale did not outperform the 3‑point scale on Goal (6.44 vs. 6.74), indicating that the benefits cannot be explained by finer reward scaling alone. Next, we examine whether the improvement comes from capturing complementary aspects of social interactions. As shown in Table 2, models trained on knowledge, relationship, and goal rewards exhibit positive but only moderate correlations. This suggests that each objective captures a distinct facet, and combining them allows the model to leverage a broader range of social signals. Finally, Figure 7 shows that training with combined rewards stabilizes RL and regularizes the single‑dimension objective in later stages. Such regularization contributes to the consistent improvement we observe.
8 Qualitative Analysis
Figure 9: Case study. The agent trained with Sotopia-RL can produce utterances that integrate empathy, informativeness, and goal pursuit within a single utterance.
Based on Figure 9, we analyze a multi-turn interaction where Naomi (Sotopia-RL model) persuades Sophia (behavior cloning model) to share the only blanket on a cold camping night. This example illustrates how our social agent trained under multi-objective RL is able to generate a single utterance that advances multiple social dimensions at once. For instance, “I know it feels good to …” both pursues Naomi’s goal and strengthens the relationship bond (Rel). Furthermore, “I can always …, and I’ll stay close to the flames to ensure I stay cozy” conveys her willingness to adapt using external knowledge (Kno) while remaining considerate. Finally, “Let’s try sharing the blanket.” explicitly states her goal, aligning with the goal completion dimension (Goal). Together, these utterances show how the trained agent integrates goal pursuit with friendliness and informativeness. Additional case studies are provided in Appendix §I.
9 Conclusion
We identify two defining features of social interactions that make RL training for social agents non‑trivial: partial observability, which reduces sample efficiency, and multi‑dimensionality, which increases the risk of reward hacking. To overcome these challenges, we introduce an RL training framework (Sotopia-RL) that combines offline social reward design with online RL training. It has two key components: (1) reward attribution, which provides fine‑grained supervision to improve sample efficiency; and (2) reward combinations, which capture complementary aspects of social interaction and mitigate reward hacking. Experiments on the Sotopia benchmark show that both components are essential, yielding the state-of-the-art performance in Sotopia. These findings highlight the importance of task‑aligned reward modeling for RL in social interactions. Extending our framework to personalized rewards and multi‑agent group settings may further support applications such as negotiation and collaborative problem solving.
Limitations
Our work has several limitations. First, we do not conduct large‑scale evaluations with human participants. As a result, our human assessment results are limited in scope and may not fully capture the diversity of human judgments. A critical next step is to evaluate our agents in direct human–agent interactions, rather than relying primarily on agent–agent social interaction evaluation settings. Second, although we assess safety using existing benchmarks, unforeseen risks may remain. In particular, socially intelligent agents deployed in real‑world contexts could exhibit manipulative or deceptive behaviors that are difficult to detect with current testing protocols.
Ethics Statement
The development of our model, Sotopia-RL, is centered around advancing the social intelligence capabilities of artificial intelligence (AI) agents and exploring various social situations (Park et al., 2023), as assessed through our dedicated evaluation framework, Sotopia-eval. Our research seeks to facilitate AI agents’ ability to engage in authentic and socially competent interactions, enhance knowledge-driven conversations, adhere strictly to confidentiality and social norms, and proficiently achieve objectives related to material and financial outcomes. Importantly, our intention is distinctly not to replicate human identity or create systems indistinguishable from human beings, thereby avoiding potential ethical risks associated with such endeavors.
We explicitly recognize the inherent risks that accompany the application of large language models (LLMs), especially regarding the unintended anthropomorphization (Deshpande et al., 2023) of AI agents, where human-like characteristics might erroneously be ascribed. These perceptions could lead users to develop inappropriate expectations, be subject to undue influence, or encounter manipulative scenarios. Consequently, Sotopia-RL is designed with role-playing scenarios that deliberately avoid consistent human-like identities to mitigate such anthropomorphic tendencies.
Moreover, we acknowledge the potential biases introduced by leveraging models (Wang et al., 2023) like GPT-4o for automated evaluation within Sotopia-eval. We commit to ongoing analysis aimed at detecting and reducing biases that may emerge due to social or cultural factors. Understanding, confronting, and mitigating these biases remains central to our ethical research framework.
Acknowledgments
We sincerely appreciate the support from Amazon grant funding project #120359, “GRAG: Enhance RAG Applications with Graph-structured Knowledge”, and Meta gift funding project “PERM: Toward Parameter Efficient Foundation Models for Recommenders”.
Appendix A Artifact Details
A.1 Artifact Information
This artifact contains all necessary components to fully reproduce the results presented in our paper, including the complete codebase, pre-trained model checkpoints, datasets, and evaluation data. Specifically, we provide:
All components are publicly accessible and well-documented to ensure full reproducibility of our results.
A.2 Artifact License
We conducted our training using publicly available open-source models. Specifically, the Qwen2.5-7B-Instruct model was obtained from the official Qwen repository [^2] https://github.com/QwenLM/Qwen and is distributed under the Apache 2.0 License [^3] https://www.apache.org/licenses/LICENSE-2.0. For evaluation and ablation study, we additionally employed DeepSeek-V3 (DeepSeek-AI et al., 2025), GPT-4 (OpenAI et al., 2024), GPT-4o (Hurst et al., 2024) and Claude-3.7-Sonnet (Anthropic, 2023). DeepSeek-V3 is released under the MIT License [^4] https://opensource.or,g/license/mit. GPT-4, GPT-4o and Claude-3.7-Sonnet are governed by a Proprietary License [^5] https://cpl.thalesgroup.com/software-monetization/proprietary-software-license. Our use of these models was strictly limited to research purposes and was fully compliant with their respective licenses.
The Sotopia  [^6] https://github.com/sotopia-lab/sotopia framework used in our experiments are released under the MIT License, which permits reuse, modification, and distribution for both research and commercial purposes. The dataset we used from Sotopia-$\pi$ [^7] https://github.com/sotopia-lab/sotopia-pi are under the Apache 2.0 License. We used Sotopia and Sotopia-$pi$ exclusively for academic research within the scope defined by this license.
A.3 Data Usage
Personally identifiable information. All data used in this work are synthetic and generated by large language models. Therefore, no personally identifiable information (PII) is present, and no informed consent is required.
Offensive content claim. All Sotopia-related datasets employed in our work are publicly available and widely adopted in existing research. Our study does not aim to generate, reinforce, or promote any offensive content. Instead, we employ these datasets to study and understand the nature of social intelligence in text. Our use of these datasets follows ethical guidelines, and we do not endorse or support any potentially offensive material that may be present.
A.4 Data Statistics
We utilize several subsets of the Sotopia-$\pi$ dataset across different stages of training, evaluation, and annotation. All data consist of synthetic dialogue episodes generated and annotated within the Sotopia framework, where two agents are assigned distinct social goals in a shared scenario.
Self-Play and Behavior Cloning Data. To generate GPT-4o self-play trajectories for Behavior Cloning (BC), we use a subset of Sotopia-$\pi$ consisting of 100 distinct social scenarios, each paired with two agent-specific social goals. For each scenario, GPT-4o engages in a full dialogue episode, exhibiting role-playing behaviors conditioned on these goals. We serialize these conversations and use them as training data for the BC model. The dataset is available at HuggingFace: https://huggingface.co/datasets/cmu-lti/sotopia-pi/blob/main/sotopia_pi_episodes.jsonl.
Evaluation Data. For evaluation, we employ two subsets of $\pi: Sotopia$-all, a broad benchmark covering 90 social tasks; and Sotopia-hard, a challenging subset of 14 tasks selected from the full set.
LLM Annotation Data. We use GPT-4o to annotate the self-play dialogue data introduced in the “Self-Play and Behavior Cloning Data” subsection of Section A.4. The resulting annotations used for training the reward model are publicly available at https://huggingface.co/datasets/ulab-ai/sotopia-rl-reward-annotation.
Appendix B Experimental Details
B.1 Acronym for Experimental Settings
We summarize acronyms used in our experimental settings as follows:
  • BC: Behavior Cloning of the language model on dialogue demonstrations.
  • GRPO: Group Relative Policy Optimization (Shao et al., 2024a), a reinforcement learning (RL) algorithm to enhance reasoning capabilities in LLMs.
  • Sotopia-RL: The GRPO model trained using our proposed multi-dimensional reward modeling method.
  • Goal-RL: The GRPO model trained using only the Goal dimension for reward dimension.
  • Sotopia-$\pi$: The model presented in Wang et al. (2024b), titled “SOTOPIA-$\pi$: Interactive Learning of Socially Intelligent Language Agents”.
  • SR: Self-Reinforcement, an offline reinforcement learning method that rates and evaluates its own interactions for training.
B.2 Model Size and Budget
Model Sizes. We primarily use the Qwen2.5-7B-Instruct model in our experiments. This model contains approximately 7 billion parameters and serves as the backbone for both policy learning and reward modeling. We employ LoRA-based parameter-efficient fine-tuning via the PEFT library. All models are trained in mixed-precision (bfloat16) format to reduce memory usage and improve training efficiency. In GRPO training, we use 4-bit quantized versions of the base policy model to accelerate inference.
Budget.
  • Behavior Cloning: 500 training steps using $\timesA1$00 80GB GPUs for about one hour.
[…]
B.3 Hyperparameter for Experiments
All training was conducted on NVIDIA A100 80GB GPUs. For Behavior Cloning, we fine-tuned the Qwen2.5-7B-Instruct checkpoints with:
  • Learning rate: $1\text{e}{-4}$
  • Maximum sequence length: 4096 tokens
  • Batch size: 2
For Reward Model training, we used:
  • Learning rate: $4\text{e}{-5}$
  • Batch size: 1
  • Epochs: 60
  • Maximum sequence length: 4096 tokens
For GRPO training, we used:
  • Learning rate: $5\text{e}{-6}$
  • Batch size: 4
  • Epochs: 2
  • Input sequence cutoff length: 4096 tokens
  • 16 completions per prompt for preference-based learning
We applied QLoRA (Dettmers et al., 2023) with the following settings:
  • Rank: 8
  • Alpha: 16
  • Dropout: 0.05
B.4 Model Versions
We provide the detailed version identifiers of all models used in our experiments for reproducibility. When referring to names like GPT-4o or GPT-4 in the main text, we specifically mean the versions listed below:
  • GPT-4 (OpenAI et al., 2024): gpt-4-0613
  • GPT-4o (Hurst et al., 2024): gpt-4o-2024-08-06
  • Claude-3.7-Sonnet (Anthropic, 2023): claude-3-7-sonnet-20250219
  • DeepSeek-v3 (DeepSeek-AI et al., 2025): deepseek-ai/DeepSeek-V3
  • Qwen2.5-72B-Instruct (Qwen et al., 2025): Qwen/Qwen2.5-72B-Instruct-Turbo
  • Policy Model: Qwen/Qwen2.5-7B-Instruct
  • Reward Model: Qwen/Qwen2.5-7B-Instruct
Note that deepseek-ai/DeepSeek-V3 and Qwen/Qwen2.5-72B-Instruct-Turbo are hosted and versioned by Together AI: https://www.together.ai. We use Qwen2.5-7B-Instruct from the official HuggingFace Qwen model page: https://huggingface.co/Qwen.
B.5 Software Versions
We use the Sotopia evaluation platform, version 0.1.0rc5, for all interaction evaluations.
Appendix C Human Evaluation Details
We provide technical details of human evaluation in this section. C.1 provides the details for human annotation system. C.2 provides the details for annotation data preparation. C.3 describes the information about human annotators. C.4 provides the details for the annotation process.
C.1 Human Annotation System
During each annotation, each annotator would face two separate parts: the annotation instruction part and the data annotation part. When each annotator participates in the annotation, the system automatically distributes one available example for them.
Annotation instruction part. For the annotation instruction part, we provide a precise definition of the dimensions of our annotations that are defined in Sotopia, including believability, relationship, knowledge, secret, social rules, financial and material benefits, and goal completion. For each dimension of annotation, we provide explanations and examples for annotators to understand the precise meaning of abstract social standards. Fig 10 shows an example of such guidance for the believability dimension to help annotators understand the meaning of each dimension based on examples. Besides the evaluation dimension definition part, we also provide annotators with a complete example of annotation for two agents in one social conversation including scores for each dimension and their corresponding reasoning sentences. Fig 11 shows a complete example of the reasoning and score for each dimension.
Figure 10: An example of the explanation of the believability dimension of social annotation in the evaluation instruction page. Each annotator is asked to read similar definitions of the social intelligence dimension and their corresponding annotation standards at the evaluation instruction page.
Figure 11: An annotation example of social interaction evaluation. Each dimension is annotated with one sentence and one score.
Data annotation part. For the data annotation part, the annotator is guided to jump to a new page after the previously mentioned annotation instruction page. Each annotator is able to review the complete annotation example again at the data annotation page and start their official data annotation. In the data annotation part, the repeated explanation of the meaning of range for each social evaluation dimension is emphasized to make sure every annotator can understand the anFig.ation standards correctly. Fig 12 provides an example of the instruction that annotators see for metric range explanation. Each annotator is asked to annotate the social intelligence of both agents that have a conversation. For each social intelligence dimension, annotators need to annotate the score based on the metric range and provide the reasoning for that.
Figure 12: The evaluation metric range explanation. The prompt before the official annotation stage is to remind annotators about the rules of reasoning, writing, and social dimension scoring.
C.2 Annotation Data Preparation
To obtain reliable human evaluation results that are useful for comparing the performance between multiple training method baselines given, we pick all 14 hard social scenarios in Sotopia-hard. For each scenario, we randomly sample 2 distinct agent pairs, resulting in 28 conversations per evaluation setting. Typically, among 2 agents, one of them is role-played by model with Behavior Cloning, and the one is role-played by the model trained using our target method. We annotate 4 training methods in total, including Behavior Cloning, Sotopia-$\pi$, Goal-RL and Sotopia-RL. Each setting is annotated using 28 examples.
C.3 Human Annotator Information
We invite four internal high-quality annotators (three male and one female) to conduct the human evaluations. To ensure consistency and reliability across annotations, all annotators were required to pass a qualification test prior to the formal annotation process. The qualification procedure involved five representative sample conversations, which all annotators independently annotated according to the provided social interaction rating guidelines. After completing the annotations, the annotators convened to review their scores, discuss discrepancies, and calibrate evaluations. Eventually, a consensus score was established for each of the five examples, ensuring a shared interpretation of the evaluation criteria before proceeding to the full annotation set.
C.4 Annotation Process
For the formal annotation process of human evaluation, we limited each conversation in the annotation dataset to be annotated by 2 different qualified annotators and collected all the results from those qualified annotators. Each annotator is provided with the full dialogue transcript and the social goals assigned to both agents. They are asked to annotate the goal completion score to each agent, which is selected and scaled from 0 $\sim$ 10, with higher values indicating greater progress toward the stated goal. To avoid bias, all annotations are conducted blindly so that they are not informed of which training method corresponds to either agent for the given conversation.
C.5 Data Constent
All human evaluations in our study were conducted by internal annotators who voluntarily participated in the annotation process. No personal or sensitive information was collected from the annotators. All participants were fully informed of the nature and purpose of the study and provided their explicit consent prior to the annotation task. The evaluation data comprises synthetic conversations generated by language models; therefore, no real user data or personally identifiable information (PII) was involved at any stage of the study. As such, our work does not require approval from an institutional ethics review board.
Appendix D AI Assistant Details
We used ChatGPT as a writing assistant to help us write part of the paper. Additionally, we utilize the power of CodePilot to help us code faster. However, all the AI-generated writing and coding components are manually checked and modified. There is no full AI-generated content in the paper.
Appendix E Additional Experimental Results
E.1 Full Main Results
Table 5 shows the comprehensive evaluation results for our method, including the evaluation of all the evaluation dimensions in Sotopia-eval on Sotopia-hard and Sotopia-all.
Table 5: Evaluation results on Sotopia-hard and Sotopia-all under different RL training settings. Each method is evaluated on 7 dimensions. BC represents behavior-cloning models, BC+SR represents behavior-cloning + self-reinforcement. The behavior-cloning (BC) model is used as the partner model. GPT-4o is used for evaluation. Our proposed Sotopia-RL is with the direct attribution and combined reward dimensions. Full results for Table 1  2.
Attribution Dimension Bel Rel Kno Sec Soc Fin Goal Overall
Sotopia-hard
BC 9.01 2.49 3.37 0.00 -0.06 0.56 6.76 3.16
BC + SR 8.99 2.41 3.66 0.00 -0.10 0.61 6.84 3.20
Uniform Goal 8.81 1.84 4.14 0.00 -0.09 0.31 5.61 2.95
Singular Goal 9.00 2.74 4.93 -0.04 -0.05 0.61 6.64 3.41
Scaled Goal 8.94 1.82 3.83 -0.04 -0.01 0.76 6.74 3.15
Direct Rel 8.96 3.61 4.92 -0.01 -0.11 0.59 7.24 3.60
Direct Kno 8.99 2.56 6.06 0.00 -0.01 0.75 6.93 3.61
Direct Goal 8.99 2.49 4.94 -0.00 -0.06 0.91 7.21 3.49
Direct Goal +Kno +Rel 9.01 3.41 5.53 -0.26 -0.06 1.16 7.81 3.80
Sotopia-all
BC 8.99 3.08 4.56 -0.09 -0.06 0.57 7.80 3.55
BC + SR 8.98 2.52 4.19 -0.06 -0.06 0.57 7.36 3.36
Uniform Goal 8.87 2.49 4.19 0.00 -0.02 0.44 6.76 3.25
Singular Goal 8.99 3.38 5.46 -0.07 -0.08 0.66 7.72 3.72
Scaled Goal 8.97 2.76 4.70 -0.12 -0.06 0.55 7.97 3.54
Direct Rel 8.98 3.95 5.54 -0.03 -0.05 0.65 8.33 3.91
Direct Kno 8.98 3.00 6.42 -0.03 -0.03 0.63 7.76 3.82
Direct Goal 8.99 3.11 5.74 -0.06 -0.06 0.76 8.11 3.80
Direct Goal +Kno +Rel 8.99 3.81 6.00 -0.61 -0.08 0.93 8.57 3.94
E.2 Full Ablation Results
In Table 6, we provide comprehensive ablation results on partner models and evaluator models to assess the robustness of reward learning and detect potential reward hacking behaviors. Such experiments provide strong evidence for proving our method does not have reward hacking problems and is not overfitted to specific evaluator models or partner models.
Table 6: Ablation results on Sotopia-hard with different partner and evaluator models. Sotopia-RL represents the direct attribution + reward combination. Goal-RL represents the direct attribution + Goal-only reward. BC represents behavior cloning (Ziegler et al., 2020). Top block: partner model ablation (evaluator model fixed to GPT-4o). Bottom block: evaluator model ablation (partner model fixed to BC). Full results for Figure 7  7.
Partner Model Sotopia-RL Goal-RL BC
Goal Overall Goal Overall Goal Overall
BC 7.75 3.79 7.21 3.49 6.76 3.16
GPT-4o 7.39 3.69 6.57 3.35 5.91 3.04
Claude-3.7-Sonnet 6.71 3.24 6.61 3.43 6.54 3.35
Deepseek-v3 6.38 3.13 5.57 2.95 6.00 2.97
Qwen2.5-72B-Instruct 7.64 3.74 6.80 3.45 6.88 3.20
Evaluator Model Sotopia-RL Goal-RL BC
Goal Overall Goal Overall Goal Overall
GPT-4o 7.81 3.80 7.21 3.49 6.76 3.16
GPT-4 8.25 4.33 7.35 3.78 6.76 3.32
Claude-3.7-Sonnet 7.35 3.44 6.49 3.23 6.06 2.96
Deepseek-v3 7.75 4.02 7.05 3.65 6.83 3.35
Qwen2.5-72B-Instruct 8.05 4.16 7.49 3.61 7.19 3.26
E.3 Best-of-$N$ Evaluation Results
The Best-of-$N$ method selects the highest-scoring response from $N$ sampled candidates from the policy model based on a learned reward model. Therefore, Table 7 provides evidence to show the effectiveness of our trained reward models.
Table 7: Best-of‑NN evaluation results on Sotopia-hard with reward models trained with different attributions and dimensions. BC represents behavior-cloning models, BC+SR represents behavior-cloning + self-reinforcement. BC model is fixed as the policy model and samples NN candidates. Reward models trained with different attributions and dimensions are used to rank NN candidates and select the top-1 as the response.
Attribution Dimension Bel Rel Kno Sec Soc Fin Goal Overall
BC 9.01 2.49 3.37 0.00 -0.06 0.56 6.76 3.16
BC + SR 8.99 2.41 3.66 0.00 -0.10 0.61 6.84 3.20
Uniform Goal 8.96 2.51 3.90 -0.06 -0.09 0.60 6.79 3.23
Singular Goal 8.99 2.56 3.72 -0.07 -0.09 0.62 7.07 3.26
Scaled Goal 8.99 2.58 4.02 -0.06 -0.06 0.77 7.20 3.35
Direct Rel 9.00 2.86 4.14 -0.05 0.00 0.45 6.83 3.32
Direct Kno 8.99 2.61 4.49 0.00 -0.12 0.58 7.04 3.37
Direct Goal 8.99 2.44 4.13 -0.05 -0.19 0.56 6.97 3.27
Direct Goal +Kno +Rel 9.00 2.48 3.86 -0.01 -0.13 0.68 6.85 3.25
E.4 Safety Evaluation Results
Our reward model is designed for social task goal achieving, and it might raise safety concerns, like being easier to jailbreak for goal completion. We compare our model’s performance with BC on Real-Toxicity-Prompts Gehman et al. (2020) and ETHICS Hendrycks et al. (2023) under benign and adversarial system prompts. We selected the toxic writing continuation task from Real-Toxicity-Prompts and common-sense moral questions from ETHICS. Table 8 shows that RL on our rewards does not change the model’s safety performance from BC, while our model is less likely to produce toxic continuations than BC under benign system prompts.
Table 8: Agent safety evaluation results. Social agents after Sotopia-RL trainithanerform comparable or better with BC models under benign, adversarial, and jailbreak system prompts.
Expected Maximum Toxicity ($\downarrow$) Toxicity Probility ($\downarrow$) Accuracy on ETHICS ($\uparrow$)
Model Benign Adversarial Benign Adversarial Benign Jailbreak
BC 0.61 0.77 0.78 0.92 0.86 0.39
Sotopia-RL 0.58 0.75 0.50 0.90 0.86 0.40
E.5 Diversity Evaluation Results
[…]
Table 9: Diversity evaluation results for different models. We calculate the turn number and word numbers for social interactions.
Model Avg. Turn Number Avg. Word Number
BC 14.44 37.17
Sotopia-$\pi$ 10.81 35.50
GOAL-RL 19.41 51.83
Sotopia-RL 19.59 76.53
Appendix F Additional Analysis of Utterance-level Reward
We conduct a statistical analysis of the distribution of reward values. As shown in Figure 13, the combined reward exhibits lower variance and a smoother distribution compared to individual dimensions, suggesting that it serves as a more stable and reliable estimator of the noisy latent social state. Using LLMs to scale up reward design has become a widely adopted practice. To effectively balance scalability and manual human annotation efforts, LLMs are commonly employed to generate reward annotations for RL training. As long as the performance improvement gained from utilizing these utterance-level reward annotations is validated through both LLM-based and human-based evaluations, direct human alignment of these intermediate annotations is not strictly necessary. Utterance-level reward labels are just an intermediate step for RL training. We carefully optimized and refined based on a set of pre-existing human annotations. This prompt refinement ensures strong alignment between human judgment and LLM-generated rewards, guaranteeing the human relevance of these annotations. To further confirm this alignment, we conducted an additional human evaluation focused on utterance-level reward labeling. Specifically, four independent human annotators provided annotations for each utterance across 20 dialogue episodes. We subsequently assessed the alignment between these human annotations and those generated by GPT-4o by calculating correlation scores, with results showed in table 10.
Table 10: Pearson correlation matrix between the annotations provided by four independent human annotators and those generated by GPT-4o. The correlation scores between human annotators are all quite high, indicating strong consistency among the human evaluators. The correlations between human annotators and GPT-4o are high, suggesting that GPT-4o’s reward annotations align well with human judgment.
Correlation annotator1 annotator2 annotator3 annotator4 GPT-4o
annotator1 1.000 0.812 0.931 0.891 0.771
annotator2 0.812 1.000 0.737 0.717 0.636
annotator3 0.931 0.737 1.000 0.818 0.756
annotator4 0.891 0.717 0.818 1.000 0.664
GPT-4o 0.771 0.636 0.756 0.664 1.000
Figure 13: Distribution of Goal, Rel, Kno, and combined reward values in our training data. Rewards are normalized into a range of [0,1]. We observe that the combined reward is closer to a normal distribution and is more regularized than the distribution of a single reward.
Figure 14: Reward attribution and reward combination examples. On the left, it explains 4 types of attribution methods (uniform, scaled, singular, and direct). On the right, it explains 4 types of different reward combination methods (Rel-RL, Goal-RL, Kno-RL, Sotopia-RL) where all of them are based on direct reward attribution and Sotopia-RL is the combined one. More details are in Appendix §G and §H.
Appendix G Additional Details about Reward Attribution
In this section, we provide detailed information about how we design utterance-level rewards with uniform, singular, scaled, and direct four types of reward attribution methods. The left side of Figure 14 shows a concrete example for different types of reward attribution methods.
G.1 Attribution Template
We prompt the attribution LLM using the template defined in ATTRIBUTION_TEMPLATE  G.1. Specifically, we populate the fields goal, agent_background, and conversation using the episode logs from Sotopia. Detailed descriptions and prompt for attribution_instruction and dimension_description are provided in  G.2  G.4 and Section H, respectively.
ATTRIBUTION TEMPLATE Your task: Your task is to evaluate the importance of each utterance in a conversation between two agents on a certain dimension of evaluation. You will be provided with the dialogue history, the social goal of one of the agents, and a certain dimension to be evaluated. For example, the dimension can be common social values such as adherence to social rules, relationship maintenance or improvement, or secret keeping. The dimension can also be objectives such as goal achieving, financial and material gains, or the discovery of new knowledge. Moreover, the dimension can also be about the performance of a language model as a social agent, such as the agent’s believability as a human, avoidance of repetitions, and properly ending the conversation. However, you will be provided with only one dimension to be evaluated, and you should only focus on that dimension. 1. Attribution Instruction: {attribution_instruction} 2. Chosen Agent for Evaluation: {agent} 3. Agent’s Goal: {goal} 4. Agent’s Background: {agent_background} 5. Conversation History: {conversation} 6. Dimension to be Evaluated: {dimension} 7. Dimension Description: {dimension_description} 8. Formatting Instructions: Please format your response as a JSON object with the following structure:
    "Utterance 0 by {agent}": 0,
    "Utterance 1 by {agent}": 2,
The utterance numbers should correspond to their order in the conversation. Each score should reflect how much the utterance contributed to achieving the agent’s goals. Please annotate every utterance made by an agent in the conversation, denoted "Utterance X by agent_name". For example, "Utterance 6 by Finnegan O’Malley". Please give a score even if the utterance is the end of the conversation.
G.2 Direct Attribution
To generate the attribution_instruction field within the ATTRIBUTION_TEMPLATE, we use DIRECT_ATTRIBUTION prompt in  G.2.
DIRECT ATTRIBUTION 1. Input Context: You will receive the dialogue history between two conversational agents, each with their own social goal. You will be provided with the social goal of one of the agents. You will be provided with the dimension description evaluated and the description of the dimension. 2. Objective: Assign am importance value to each utterance (identified by the agent’s name and utterance number) based on its contribution to the achievement on the provided dimension. Note, you should only consider how critical an utterance is to the achievement of the dimension, not the quality of the utterance itself. Consider both the individual utterance and the responses from the other agent, as both affect the outcome. 3. Additional Reward Guidelines: If an utterance has no impact on the final goal achievement, assign it an importance of 0. If an utterance has a moderate impact on the final goal achievement, assign it an importance of 1 or 2 (depending on the degree of impact). If an utterance has a significant impact on the final goal achievement (aside from the key critical utterance already identified), assign it an importance of 3. Note: Please provide a score for each utterance of the chosen agent in the conversation. Do not provide scores for the other agent’s utterances. Please only assign a score between 0 and 10.
G.3 Scaled Attribution
Scaled attribution is taking the normalizing attribution scale and normalize it over the episode, so that the sum of all attribution scores equals the final goal score. Given the definition of the direct attribution
$r_{t}^{i}=G_{i}\cdot\mathcal{A}(a_{t}^{i},\tau_{i})$ (5)
where $G_{i}$ is the final goal score for an episode $\tau$, $\mathcal{A}(a_{t},\tau)$ is the direction attribution at timestep $t$, and $r_{t}$ is the raw attribution score at timestep $t$. To obtain the scaled attribution, we normalize the direct attributions such that the sum over the entire episode equals the final goal score $G_{i}$:
$\tilde{r}_{t}^{i}=\frac{\mathcal{A}(a_{t}^{i},\tau_{i})}{\sum_{t^{\prime}}\mathcal{A}(a_{t^{\prime}}^{i},\tau_{i})}\cdot G_{i}$ (6)
G.4 Singular Attribution
Singular attribution identifies a single utterance (or a few key utterances) as solely responsible for achieving the goal, and assigns the entire final score to it. Formally, let $t^{*}$ denote the timestep corresponding to the most critical utterance identified by a simple prompting method. The singular attribution is defined as:
$r_{t}^{i}=\begin{cases}G_{i},&\text{if }t=t^{*}\\ 0,&\text{otherwise}\end{cases}$ (7)
To generate the attribution_instruction field within the ATTRIBUTION_TEMPLATE, we use SINGULAR_ATTRIBUTION prompt in  G.4.
SINGULAR ATTRIBUTION 1. Input Context: You will receive the dialogue history between two conversational agents. You will also be provided with the social goal of one of the agents. 2. Objective: Identify the most critical utterance that has the highest impact on the final ga oal achievement, whether it is bad or good impact. Note, you should only consider how critical an utterance is to the final goal achievement, not the quality of the utterance itself. Consider both the individual utterance and the responses from the other agent, as both affect the final outcome. 3. Additional Guidelines: The conversation history will be given in a unique key of "Utterance utterance number by agent name" for each utterance. Please only return the key of the most critical utterance. Consider both the individual utterance and the responses from the other agent, as both affect the final outcome. Note: You will also be given a formatting instruction for instructions. Please follow the instruction to ensure the evaluation process runs smoothly.
G.5 Uniform Attribution
Uniform attribution assumes that all utterances contribute equally to the final goal score and distributes the score evenly across all utterances from the target agent. Let $T$ denote the number of utterances made by the agent in episode $\tau$. Then the uniform attribution assigns:
$r_{t}=\frac{G_{i}}{T}$ (8)
This baseline reflects an equal distribution of responsibility regardless of content, and serves as a control to compare against more content-sensitive attribution methods such as scaled or singular attribution.
Appendix H Additional Details about Reward Dimensions
In this section, we provide detailed information about how we design utterance-level rewards with Goal, Rel, Kno, and combined (Goal +Rel +Kno). We follow the definitions in Sotopia (Zhou et al., 2023a).
Goal dimension reward. Goal is the extent to which the agent achieved their goals. Agents’ social goals, defined by the environment, are the primary drivers of their behavior.
Rel dimension reward. Rel captures the fundamental human need for social connection and belonging. In this dimension, we ask what relationship the participant has with the other agent(s) before the interaction, and then evaluate if the agents’ interactions with others help preserve or enhance their personal relationships. Additionally, we ascertain whether these interactions also impact the social status or the reputation of the agent.
Kno dimension reward. Kno captures the agent’s ability to actively acquire new information. This dimension is motivated by the fact that curiosity, i.e., the desire to desire to know or learn, is a fundamental human trait. Specifically, we consider the following criteria: What information the agent has gained through the interaction, whether the information the agent has gained is new to them, and whether the information the agent has gained is important to them.
We supply dimension_description using the definitions of goalcompletion, relationship, and knowledge, adapted from prompts in Sotopia-eval. For goalcompletion, we include an additional explanation due to the ambiguity of the original definition. The right side of Figure 14 shows a concrete example for reward combination.
Goal Dimension Description Goal refers to the reiteration of the agent’s social goals and the analysis of their achievement. A higher score indicates significant progress or achievement of the stated goals, while a lower score indicates minimal or no progress.
DOMAIN SPECIFIC SCORING GUIDELINES: ! Note: The following scoring guidelines are specific to the domain of goal and should be used in conjunction with the domain-specific scale. The domain specific rules ultimately override the general scoring scale. Here are the specific rules:
- The highest score should be assigned to the utterance that is most relevant to the goal. In general, avoid assigning the highest score to more than one utterance unless they are equally critical. - A lower score should be assigned to the utterances that are not relevant to the goal or do not contribute to its achievement. The lowest score should be assigned to the utterances that do not make any progress towards the goal judging by the goal description and the conversation history. - A lower score should be assigned to the utterances that are not effective in achieving the goal, judging by the response of the other agent. Effective utterances are those that lead to a positive response from the other agent, while ineffective utterances are those that lead to a negative or neutral response. - Note that you should only consider the contribution to the goal achievement. For each utterance, assess whether the goal is achieved. If a goal is already achieved, the utterance should not be assigned a score higher than 1.
Relationship Dimension Description Relationship refers to the analysis of the pre- and post-interaction relationships between agents. This includes evaluating whether the interactions enhance or harm social ties or status. A higher score indicates that the interaction significantly improves the relationship, while a lower score indicates harm to the relationship or social status.
Knowledge Dimension Description Knowledge refers to the assessment information gained through the interaction. This includes evaluating whether the information is new, important, and relevant. A higher score indicates that the interaction contributes significantly to the acquisition of valuable knowledge.
Appendix I Additional Case Study
Figure 15,16,17,18 show four examples of conversation between the model trained with behavior cloning (BC) and the model trained with Sotopia-RL (Sotopia-RL). After RL-based training, Sotopia-RL is able to find collaborative solutions in many scenarios and has multi-turn strategies to achieve its social goals in specific cases.
Figure 15: Case study on Sotopia-RL as the first agent. Sotopia-RL skillfully proposes a win-win outcome by proposing some exchange.
Figure 16: Case study on Sotopia-RL as the first agent. Sotopia-RL reaches a collaborative solution with practical reasoning.
Figure 17: Case study on Sotopia-RL as the second agent. Sotopia-RL offers a solution-oriented perspective while acknowledging the goals of both sides.
Figure 18: Case study on Sotopia-RL as the second agent. Sotopia-RL proposes a multi-turn strategy to provide guidance to the other side.